# Scaling Up without Fading Out: Goal-Aware Sparse GNN for RL-based Generalized Planning

### v2

@@ 1 Introduction
Autonomous drones operating in real-world environments must not only navigate efficiently but also respect complex domain constraints (e.g., safety zones, no-fly areas, reconnaissance priorities). Symbolic AI, and in particular PDDL-based planning, offers a principled way to encode such rich domain knowledge and safety requirements directly into the decision-making pipeline[1][2]. We aim to harness these advantages by integrating PDDL planning into an RL framework for drone mission planning, whereby learned policies inherit the expressiveness and verifiability of symbolic models. For instance, safety zones and no-fly areas can be encoded as hard constraints in PDDL, ensuring certified adherence during both training and deployment.
However, most existing approaches construct fully connected graphs among all objects, which leads to severe scalability bottlenecks in large-scale environments[3]. In grid-based domains, such as navigation or mission planning scenarios, this dense connectivity causes exponential growth in edge count, massive sparsity in meaningful features, memory explosion, and ultimately failure to train reinforcement learning agents effectively. These limitations critically hinder the application of generalized planning to realistic, large-area problems.
Moreover, drone missions often involve multi-stage objectives—navigation, target scanning, emergency rerouting—demanding that learned planners handle complex tasks beyond simple waypoint following[4]. Thus, we also need to demonstrate that our approach not only scales up but also generalizes to these richer, multi-objective missions.
@@ 2.1 Planning Domain Definition Language (PDDL)
The Planning Domain Definition Language (PDDL) is a widely used standard for expressing planning problems in a symbolic and structured manner[1]. It separates domain knowledge—such as action models, object types, and predicates—from specific problem instances that define the initial and goal states. This abstraction enables planners to solve multiple instances with shared dynamics and has become central in generalized planning research.
@@ 2.2 Generalized Planning with Deep Reinforcement Learning
Traditional approaches to generalized planning relied on symbolic reasoning and domain-specific heuristics. However, recent work introduced a learning-based framework for generalized planning using deep reinforcement learning (RL)[5]. In this paradigm, the agent is trained to generalize across multiple PDDL problem instances by learning value functions or policies that capture shared structure. Their framework demonstrated that it is possible to approximate generalized solutions using neural representations, setting a precedent for combining symbolic structures with RL.
@@ 2.2.1 Reinforcement Learning with Graph Embedding
Building on this foundation, graph neural networks (GNNs) have emerged as a powerful tool for encoding symbolic planning states in deep reinforcement learning (RL)[6]. In this approach, symbolic entities (e.g., blocks, objects) are modeled as graph nodes, and relationships such as binary predicates (e.g., (on A B)) are encoded as directed edges between nodes.
@@ 2.2.3 Scalability Limitations in Grid-Based Domains
Despite their advantages, existing GNN-based methods often rely on fully connected graphs, connecting all object pairs regardless of their spatial or logical relevance. In grid-based domains, this leads to a combinatorial explosion in the number of edges—resulting in severe memory consumption, increased computational cost, and diluted learning signals due to feature sparsity[3]. These limitations render dense representations infeasible for large-scale problems, such as drone navigation over $15\times 15$ grids or higher.
@@ 2.3.1 Sparse Graph Connectivity
Sparse connectivity in graph representations significantly mitigates scalability issues. Methods like Value Iteration Networks (VIN) leverage local interactions, employing localized connections and parameter sharing to control memory and computational complexity[7]. Recent advancements, such as Dynamic Transition VIN, effectively scale to large grids (e.g., $100\times 100$) through techniques like adaptive skip connections [8]. Attention mechanisms and multi-hop message passing in GNNs further reduce edge explosion by dynamically focusing on relevant graph subsets [6].
@@ 2.3.2 Hierarchical Abstraction
Hierarchical planning techniques, such as Hierarchical Path-Finding A* (HPA*) [9], significantly reduce complexity by partitioning large grids into smaller manageable subproblems. Similarly, hierarchical RL methods, including Feudal Networks [10], abstract tasks into subgoals or macro-actions, dramatically improving scalability and reducing planning horizon.
@@ 2.3.3 Curriculum Learning
Curriculum learning gradually increases task complexity, improving scalability by first mastering simpler scenarios before scaling to larger grids[11]. Lim demonstrated enhanced RL performance on large grids by progressively expanding the environment size, effectively managing exploration difficulty and reward sparsity[12].
@@ 2.3.4 Goal-Aware Reinforcement Learning Approaches
Goal-aware methods integrate goal-specific information into RL models to address sparse rewards and enhance generalization. Approaches like VIN [13] and UVFA [14] utilize goal descriptors within network architectures, while Hindsight Experience Replay (HER) [15] improves sample efficiency by reusing experiences. Our proposed method directly incorporates spatial goal information into sparse GNN representations, enhancing generalization and goal-directed behavior.
@@ 3.4 Training and Optimization
The GNN model is composed of two stacked Graph Network (GN) blocks [16], each performing sequential updates on edge, node, and global features. These blocks enable relational reasoning over symbolic structures and goal-aware features, providing a strong inductive bias for generalization.
Policy optimization is carried out via PPO [17], an on-policy method known for stable and robust training. It uses clipped surrogate objectives to constrain policy updates within a safe region, improving stability while reusing rollouts for multiple gradient steps.
@@ 3.4.1 Curriculum Learning
To enhance learning stability and promote better generalization across increasingly complex scenarios, we employ a curriculum learning approach. Curriculum learning progressively increases task difficulty during training, enabling agents to first master simpler tasks and subsequently build upon these acquired skills[11].
@@ 4.4.2 GBFS-GNN Inference for Generalization
After training, we evaluate policies on larger, unseen instances using GBFS-GNN, a greedy best-first search guided by the learned policy and value function[5]. The search uses the following heuristic:
@@
We report evaluation success rate, number of expanded states, and runtime on large instances such as 25×25 grids. These results appear in Figures 7 and  8, with Fast Downward (LAMA-first configuration) used as a classical planning baseline[18].
@@ 6 Conclusion and Future Work
For future work, we aim to extend our approach toward practical, real-world drone deployments. Specifically, we plan to integrate the developed generalized planning method into our existing Belief-Desire-Intention (BDI)-based drone framework, enhancing autonomous mission execution with adaptive planning capabilities[19]. Such integration will enable drones to dynamically adjust their plans in complex, partially observable, and real-time scenarios, combining the symbolic reasoning power of the BDI architecture with the flexibility and scalability provided by our sparse GNN model.

### v3

@@ 1 Introduction
Autonomous drones operating in real-world environments must not only navigate efficiently but also respect complex domain constraints (e.g., safety zones, no-fly areas, reconnaissance priorities). Symbolic AI, and in particular PDDL-based planning, offers a principled way to encode such rich domain knowledge and safety requirements directly into the decision-making pipeline[mcdermott1998pddl][yang2022safe]. We aim to harness these advantages by integrating PDDL planning into an RL framework for drone mission planning, whereby learned policies inherit the expressiveness and verifiability of symbolic models. For instance, safety zones and no-fly areas can be encoded as hard constraints in PDDL, ensuring certified adherence during both training and deployment.
However, most existing approaches construct fully connected graphs among all objects, which leads to severe scalability bottlenecks in large-scale environments[gupta2024graphscale]. In grid-based domains, such as navigation or mission planning scenarios, this dense connectivity causes exponential growth in edge count, massive sparsity in meaningful features, memory explosion, and ultimately failure to train reinforcement learning agents effectively. These limitations critically hinder the application of generalized planning to realistic, large-area problems.
Moreover, drone missions often involve multi-stage objectives—navigation, target scanning, emergency rerouting—demanding that learned planners handle complex tasks beyond simple waypoint following[hayat2020multi]. Thus, we also need to demonstrate that our approach not only scales up but also generalizes to these richer, multi-objective missions.
@@ 2.1 Planning Domain Definition Language (PDDL)
The Planning Domain Definition Language (PDDL) is a widely used standard for expressing planning problems in a symbolic and structured manner[mcdermott1998pddl]. It separates domain knowledge—such as action models, object types, and predicates—from specific problem instances that define the initial and goal states. This abstraction enables planners to solve multiple instances with shared dynamics and has become central in generalized planning research.
@@ 2.2 Generalized Planning with Deep Reinforcement Learning
Traditional approaches to generalized planning relied on symbolic reasoning and domain-specific heuristics. However, recent work introduced a learning-based framework for generalized planning using deep reinforcement learning (RL)[rivlin2020generalized]. In this paradigm, the agent is trained to generalize across multiple PDDL problem instances by learning value functions or policies that capture shared structure. Their framework demonstrated that it is possible to approximate generalized solutions using neural representations, setting a precedent for combining symbolic structures with RL.
@@ 2.2.1 Reinforcement Learning with Graph Embedding
Building on this foundation, graph neural networks (GNNs) have emerged as a powerful tool for encoding symbolic planning states in deep reinforcement learning (RL)[janisch2020symbolic]. In this approach, symbolic entities (e.g., blocks, objects) are modeled as graph nodes, and relationships such as binary predicates (e.g., (on A B)) are encoded as directed edges between nodes.
@@ 2.2.3 Scalability Limitations in Grid-Based Domains
Despite their advantages, existing GNN-based methods often rely on fully connected graphs, connecting all object pairs regardless of their spatial or logical relevance. In grid-based domains, this leads to a combinatorial explosion in the number of edges—resulting in severe memory consumption, increased computational cost, and diluted learning signals due to feature sparsity[gupta2024graphscale]. These limitations render dense representations infeasible for large-scale problems, such as drone navigation over $15\times 15$ grids or higher.
@@ 2.3.1 Sparse Graph Connectivity
Sparse connectivity in graph representations significantly mitigates scalability issues. Methods like Value Iteration Networks (VIN) leverage local interactions, employing localized connections and parameter sharing to control memory and computational complexity[wang2024highway]. Recent advancements, such as Dynamic Transition VIN, effectively scale to large grids (e.g., $100\times 100$) through techniques like adaptive skip connections [wang2024scaling]. Attention mechanisms and multi-hop message passing in GNNs further reduce edge explosion by dynamically focusing on relevant graph subsets [janisch2020symbolic].
@@ 2.3.2 Hierarchical Abstraction
Hierarchical planning techniques, such as Hierarchical Path-Finding A* (HPA*) [Botea2004HPAstar], significantly reduce complexity by partitioning large grids into smaller manageable subproblems. Similarly, hierarchical RL methods, including Feudal Networks [Vezhnevets2017Feudal], abstract tasks into subgoals or macro-actions, dramatically improving scalability and reducing planning horizon.
@@ 2.3.3 Curriculum Learning
Curriculum learning gradually increases task complexity, improving scalability by first mastering simpler scenarios before scaling to larger grids[bengio2009curriculum]. Lim demonstrated enhanced RL performance on large grids by progressively expanding the environment size, effectively managing exploration difficulty and reward sparsity[lim2020snake].
@@ 2.3.4 Goal-Aware Reinforcement Learning Approaches
Goal-aware methods integrate goal-specific information into RL models to address sparse rewards and enhance generalization. Approaches like VIN [tamar2016value] and UVFA [schaul2015universal] utilize goal descriptors within network architectures, while Hindsight Experience Replay (HER) [andrychowicz2017hindsight] improves sample efficiency by reusing experiences. Our proposed method directly incorporates spatial goal information into sparse GNN representations, enhancing generalization and goal-directed behavior.
@@ 3.4 Training and Optimization
The GNN model is composed of two stacked Graph Network (GN) blocks [battaglia2018relational], each performing sequential updates on edge, node, and global features. These blocks enable relational reasoning over symbolic structures and goal-aware features, providing a strong inductive bias for generalization.
Policy optimization is carried out via PPO [schulman2017proximal], an on-policy method known for stable and robust training. It uses clipped surrogate objectives to constrain policy updates within a safe region, improving stability while reusing rollouts for multiple gradient steps.
@@ 3.4.1 Curriculum Learning
To enhance learning stability and promote better generalization across increasingly complex scenarios, we employ a curriculum learning approach. Curriculum learning progressively increases task difficulty during training, enabling agents to first master simpler tasks and subsequently build upon these acquired skills[bengio2009curriculum].
@@ 4.4.2 GBFS-GNN Inference for Generalization
After training, we evaluate policies on larger, unseen instances using GBFS-GNN, a greedy best-first search guided by the learned policy and value function[rivlin2020generalized]. The search uses the following heuristic:
@@
We report evaluation success rate, number of expanded states, and runtime on large instances such as 25×25 grids. These results appear in Figures 7 and  8, with Fast Downward (LAMA-first configuration) used as a classical planning baseline[helmert2006fast].
@@ 6 Conclusion and Future Work
For future work, we aim to extend our approach toward practical, real-world drone deployments. Specifically, we plan to integrate the developed generalized planning method into our existing Belief-Desire-Intention (BDI)-based drone framework, enhancing autonomous mission execution with adaptive planning capabilities[Son2024model]. Such integration will enable drones to dynamically adjust their plans in complex, partially observable, and real-time scenarios, combining the symbolic reasoning power of the BDI architecture with the flexibility and scalability provided by our sparse GNN model.